a.) A 2.50 L flask was used to collect a 2.65 g sample of propane gas, . After the sample was collected, the gas pressure was found to be 743 mmHg. What was the temperature of the propane in the flask? ______°C

b.)Butane,C4H10 , is an easily liquefied gaseous fuel. Calculate the density of butane gas at 0.791 atm and 24°C. Give the answer in grams per liter.

Density = ______g/L

c.) A 0.488-g sample of a colorless liquid was vaporized in an evacuated 250.-mL flask at 121°C to give a pressure of 786 mmHg. What is the molecular weight of this substance?

Molecular weight = ______amu

d.)

You fill a balloon with helium gas to a volume of 4.68 L at 23°C and 789 mmHg. What would the volume of helium be if its pressure changed to 632 mmHg but the temperature was unchanged?

_______L
Answer #1

Solution a)

Mass of propane= 2.65 g

Pressure = 743 mm Hg

Let's convert pressure mm Hg into atm.

P = (743/760) atm

= 0.977 atm

Volume = 2.5 L

Mass of propane= 2.65 g

Molar Mass of Propane = C3H8 = 44 g/mol

So, Number of moles of Propane = ( 2.65g ÷ 44 g/mol )

= 0.06 mol

Now, According to the ideal gas,

PV = nRT

or, T = PV / nR

= (0.977 atm X 2.5 L) ÷ ( 0.06 mol X 0.0821 L atm mol-1 K-1)

= 495 K

= (495- 273) oC

= 222 oC